OpenStudy (anonymous):

x + 2y = 8 x + 2y = -4 What is the solution

OpenStudy (anonymous):

x + 2y = 8 =x=8-2y

OpenStudy (anonymous):

x + 2y = -4 =x=-2y-4

OpenStudy (anonymous):

no solution, this system represent two parallel lines meaning they don't intersect so there is no solution.